# Graph-Reward-SQL: Execution-Free Reinforcement Learning for Text-to-SQL via Graph Matching and Stepwise Reward

### v2

@@ 4.3 ROT/RelNode Partial Matching (RelPM)
Similar to ASTs, RelNode can also be used to evaluate SQL similarity through graph matching. RelPM (Zhan et al., 2025) is a rule-based matching algorithm that assesses the similarity of SQLs based on their RelNode representations, denoted $\mathcal{G}_{\hat{q}}$ and $\mathcal{G}_{q^{\star}}$, respectively. A comparable approach, applied to AST structures, is known as AstPM (Zhan et al., 2025). Both algorithms adopt a hierarchical partial matching strategy and derive a global similarity score based on the Precision and Recall of node-level matching results. At the node level, matches are determined by comparing each generated node $n^{\prime}\in\mathcal{G}_{\hat{q}}$ with all the candidate nodes $n\in\mathcal{G}_{q^{\star}}$ in the reference tree. A match is established when two nodes have the same operator type and value. Additionally, a matching score is computed by comparing their subgraphs, and the candidate node with the highest matching score is selected as the final match. Further details are provided in Appendix L.
@@ 4.4 Reward Function Design
($\mathcal{R}&{}_{t}(\hat{q},q^{*})=\;\mathbbm{1}(\textit{cond}_{\text{eos}})\cdot\left[R_{\text{GMNScore}}(\hat{q},q^{*})-\beta R_{\text{kl}}(\hat{q}_{<t})\right]\\ &+\mathbbm{1}(\textit{cond}_{\text{sub}})\cdot\left[R_{\text{StepRTM}}(\hat{q}_{\leq t},q^{*})-\beta R_{\text{kl}}(\hat{q}_{<t})\right]\\ &+\mathbbm{1}(\neg\textit{cond}_{\text{eos}})\cdot\mathbbm{1}(\neg\textit{cond}_{\text{sub}})\cdot\left[-\beta R_{\text{kl}}(\hat{q}_{<t})\right],\end{aligned}$
@@ 4.5 Outcome Reward: GMNScore
$R_{\text{GMNScore}}(\hat{q},q^{\star})=\begin{cases}-1,&\text{if syntax error}\\ -0.6,&\text{if ROT error}\\ \max(0,\mathcal{M}_{\text{GMN}}+1)&\end{cases}\end{aligned}$
The GMNScore formulation introduces graded penalties for SQL queries that trigger syntax errors or ROT parsing errors [^1] Refers to failures in converting SQL into a ROT using Apache Calcite. For example, the query WITH sub1 AS (SELECT id, name FROM users) SELECT age FROM sub1; leads to an error “Column ‘age‘ not found in any table”.. For all other cases, we rescale the FuncEvalGMN similarity score $\mathcal{M}_{\text{GMN}}$ (which lies in the range $(-\infty,0]$) to the interval $[0,1)$ by first applying an affine shift and then rectifying any negative values to zero.
@@ 4.6 Stepwise Reward: StepRTM
Current ETL (Extract, Transform, Load) pipelines rarely execute their logic in a single step. Instead, analysts break the workflow into a detailed plan of subqueries, where each subquery progressively transforms the data until the query is complete.CTEs are the standard method for expressing this plan:
@@
In most cases, CTEs enhance the readability of complex SQL by providing clear representations of intermediate steps in an ETL pipeline. These steps not only facilitate data transformation but also offer a natural way of evaluating in a stepwise way.
@@
where $\mathcal{M}_{i}=\bigcup_{j=1}^{i-1}\mathcal{G}_{j}$ represents all the matched subgraphs parsed from the first $i$ subqueries, $\mathcal{G}_{j}$ denotes the maximal matched subgraph in the reference query that aligns with the $i$-th subquery $\hat{q}_{i}$. This formulation prevents repeated rewards for the same reference node and ensures that the overall signal reflects the incremental semantic coverage of the target query. This stepwise supervision improves training performance by providing richer intermediate feedback, facilitating the generation of correct SQL queries.
@@ 5 Experimental Setup
Our experiments are primarily conducted on the Spider and BIRD benchmarks. The Spider dataset (Yu et al., 2018) contains 10,181 natural language questions paired with 5,693 complex SQL queries across 138 domains. The BIRD dataset (Li et al., 2024b) consists of 12,751 questions spanning more than 37 professional fields. We use the training split of the Spider dataset for training and the development splits of both for evaluation. Additionally, the 200k-Text2SQL dataset is used for warmup before the PPO. Further details about the datasets are provided in Appendix A.
@@
We compare our proposed reward models with several representative baselines. First, we use EX, a widely adopted reward signal in recent studies (Nguyen et al., 2025; Berdnyk and Collery, 2025; Ma et al., 2025; Pourreza et al., 2025b). To evaluate the efficacy of model-based reward mechanisms, we include the LLM-based BTRM (Christiano et al., 2017; Zeng et al., 2025b), trained using the DeepSeek-Coder-1.3B-Ins as the backbone, as detailed in Appendix K. Additionally, we incorporate AstPM and RelPM (Zhan et al., 2025) as matching-based reward model baselines, following recent work (Shojaee et al., 2023; Nguyen et al., 2025; Pourreza et al., 2025b).
@@
Following Gao et al. (2024); Li et al. (2024a); Yang et al. (2024), we use the Test Suite (TS) (Zhong et al., 2020) as the evaluation metric. TS assesses correctness across multiple augmented databases, providing a more robust evaluation. Further details are provided in Appendix B.
@@
Prior to PPO training, we performed supervised fine-tuning (SFT) using two cold-start datasets. First, we sampled a subset from the 200k-Text2SQL dataset, matching the size of the Spider training set, and trained the Deepseek-Coder-1.3B/6.7B-ins for two epochs. To promote the generation of CTE SQL queries in the stepwise reward PPO experiments, we converted BIRD data into CTE format to prepare a warm-up dataset referred to as CTE-SFT. Additional details about hyperparameter are provided in Appendix C.
@@ 6.1 Reward Performance Comparison
GMNScore can replace the EX, thereby eliminating dependence on SQL execution and database environments. As demonstrated in Table 8, GMNScore achieves the highest average TS for the 1.3B and 6.7B models, highlighting the importance of well-designed reward signals in RL. Another notable observation is that RelPM outperforms AstPM, with improvements of 2.53% and 1.71% for the two model sizes. The better performance of the former over the latter can be attributed to the use of normalized logical plans for SQL parsing in ROT, which are less susceptible to surface-level syntactic differences. This also provides an effective representation for two proposed reward models. GMNScore learns deep semantic information via graph-level embeddings, bypassing the need for execution-result comparisons and thus mitigating false-positive noise. Additionally, GMNScore eliminates the necessity of constructing and maintaining databases, offering a lightweight solution for large-scale Text-to-SQL RL. Case studies are provided in Appendix  Q.
@@
DeepSeek-Coder-1.3B-Ins 39.56 11.34 25.45
+ SFT 57.74 13.30 35.52
+ CTE-SFT 46.81 14.02 30.42
+ PPO w/ RelPM & SFT 62.86 14.67 38.77
+ PPO w/ RelPM & CTE-SFT 62.57$\downarrow$0.29 19.36$\uparrow$4.69 40.97$\uparrow$2.2
+ PPO w/ RelPM & StepRTM & CTE-SFT 64.31$\uparrow$1.74 19.43$\uparrow$0.07 41.87$\uparrow$0.90
+ PPO w/ EX & SFT 65.28 17.21 41.25
+ PPO w/ EX & CTE-SFT 65.09$\downarrow$0.19 18.97$\uparrow$1.76 42.03$\uparrow$0.78
+ PPO w/ EX & StepRTM & CTE-SFT 67.89$\uparrow$2.80 19.75$\uparrow$0.78 43.82$\uparrow$1.79
+ PPO w/ GMNScore & SFT 67.70 16.10 41.90
+ PPO w/ GMNScore & CTE-SFT 68.57$\uparrow$0.87 20.40$\uparrow$4.30 44.49$\uparrow$2.59
+ PPO w/ GMNScore & StepRTM & CTE-SFT 68.67$\uparrow$0.10 21.97$\uparrow$1.57 45.32$\uparrow$0.83
Table 3: TS Performance of the Deepseek-Coder-1.3B-Ins model trained with the integration of CTE-SFT warmup and StepRTM, which consistently improves performance. ( indicating the use of a warmup phase.)
The integration of StepRTM as a stepwise reward further enhances performance. As shown in Table 3, combining CTE-SFT with StepRTM consistently results in performance improvements across various outcome reward models. Notably, our framework, which integrates GMNScore alongside StepRTM, achieves the highest overall performance. Specifically, we observe a 5.87% improvement on the BIRD dataset and a 0.97% increase on the Spider dataset. These results suggest that the BIRD dataset, which is inherently more challenging due to its diverse database and query complexity, benefits more significantly from our proposed stepwise reward.
@@ 6.2 Effectiveness of GMNScore with GRPO
Our proposed GMNScore is effective not only with PPO but also when applied to GRPO. We trained the Qwen2.5-Coder-7B/14B-Ins with PPO and GRPO. As shown in Figure 3, the results consistently demonstrate that GMNScore outperforms EX in these two RL protocols, underscoring its robustness and effectiveness.
6.3 Case Study: CTE SQL with StepRTM
@@
Table 4 presents two cases that demonstrate how the stepwise reward model enhances both correctness and structural clarity. Each case compares the reference SQL, a failed SQL query generated by a model trained solely with an outcome-based reward, and the CTE SQL query generated by a model trained with StepRTM. In the first case, the failed SQL incorrectly retrieves data from the comments table instead of the intended posts table. The CTE SQL resolves this by decomposing the task into clear subqueries: first locating the target user, then aggregating the scores of that user’s posts. In the second case, the failed SQL hardcodes gender identifiers, leading to errors in filtering. In contrast, the CTE SQL uses two dedicated subqueries to correctly filter for male superheroes and extract their superpowers.
6.4 Analysis of GMNScore Accuracy
Figure 4: AUC between reward scores and execution result during training. GMNScore exhibits superior consistency, achieving a rate of over 97.6%.
Experimental results demonstrate the effectiveness of GMNScore as a reward model in PPO, significantly outperforming BTRM. We analyze the correlation between these two reward signals and actual execution outcomes during PPO training. As shown in Figure 4, GMNScore consistently maintains a high correlation [^2] AUC is used as a metric to assess accuracy Zhan et al. (2025). In this context, 100% does not represent the best performance, as false negatives are present in the execution outcomes. with the execution results. This indicates that GMNScore provides a more stable and precise reward signal than BTRM during training, contributing to its superior performance.
@@ 7 Discussion
The GMNScore introduced in this paper offers an alternative to EX while remaining fully compatible with other reward models. As detailed in Appendix G, we extend our investigation beyond the StepRTM integration (Section 6.1) by applying hybrid outcome reward models, which further improve performance. This finding is consistent with previous work using multiple outcome reward models (Pourreza et al., 2025b; Ma et al., 2025).
We chose to omit the hints (e.g. age = year - birth_year) available in the BIRD dataset, as these are absent in most Text-to-SQL datasets, such as Spider. As a result, the performance results reported in our study are lower than those found in works that utilized the BIRD dataset’s hints.
@@ 8 Conclusion
We propose Graph-Reward-SQL, a RL framework for Text-to-SQL that replaces execution-based rewards with the outcome reward model GMNScore and the stepwise reward model StepRTM. By eliminating the dependency on execution during training, it significantly improve training efficiency. Extensive experiments demonstrate our proposed reward models achieve superior performance. It paves the way for more scalable and efficient RL-based Text-to-SQL.

### v3

@@ 4.3 ROT/RelNode Partial Matching (RelPM)
Similar to AST, RelNode can also be used to evaluate SQL similarity through graph matching. RelPM (Zhan et al., 2025) is a rule-based matching algorithm that assesses the similarity of SQLs based on their RelNode representations, denoted $\mathcal{G}_{\hat{q}}$ and $\mathcal{G}_{q^{\star}}$, respectively. A comparable approach, applied to AST structures, is known as AstPM (Zhan et al., 2025). Both algorithms adopt a hierarchical partial matching strategy and derive a global similarity score based on the Precision and Recall of node-level matching results. At the node level, matches are determined by comparing each generated node $n^{\prime}\in\mathcal{G}_{\hat{q}}$ with all the candidate nodes $n\in\mathcal{G}_{q^{\star}}$ in the reference tree. A match is established when two nodes have the same operator type and value. Additionally, a matching score is computed by comparing their subgraphs, and the candidate node with the highest matching score is selected as the final match. Further details are provided in Appendix N.
@@ 4.4 Reward Function Design
$\begin{aligned} \mathcal{R}&{}_{t}(\hat{q},q^{*})=\;\mathbbm{1}(\textit{cond}_{\text{eos}})\cdot\left[R_{\text{GMNScore}}(\hat{q},q^{*})-\beta R_{\text{kl}}(\hat{q}_{<t})\right]\\ &+\mathbbm{1}(\textit{cond}_{\text{sub}})\cdot\left[R_{\text{StepRTM}}(\hat{q}_{\leq t},q^{*})-\beta R_{\text{kl}}(\hat{q}_{<t})\right]\\ &+\mathbbm{1}(\neg\textit{cond}_{\text{eos}})\cdot\mathbbm{1}(\neg\textit{cond}_{\text{sub}})\cdot\left[-\beta R_{\text{kl}}(\hat{q}_{<t})\right],\end{aligned}$
@@ 4.5 Outcome Reward: GMNScore
[…]
The GMNScore formulation introduces graded penalties for SQL queries that trigger syntax errors or ROT parsing errors [^1] Refers to failures in converting SQL into a ROT using Apache Calcite. For example, the query WITH sub1 AS (SELECT id, name FROM users) SELECT age FROM sub1; leads to an error “Column ‘age‘ not found in any table”.. For all other cases, we rescale the similarity metric $\mathcal{M}_{\text{GMN}}$ (which lies in the range $(-\infty,0]$) to the interval $[0,1)$ by first applying an affine shift and then rectifying any negative values to zero.
@@ 4.6 Stepwise Reward: StepRTM
Current ETL (Extract, Transform, Load) pipelines rarely execute their logic in a single step. Instead, analysts break the workflow into a detailed plan of subqueries, where each subquery progressively transforms the data until the query is complete. This logic is typically expressed using CTEs, as demonstrated by the simplified example below:
@@
In most cases, CTEs enhance the readability of complex SQL by providing clear representations of intermediate steps in an ETL pipeline. These steps not only facilitate data transformation but also offer a natural way to evaluate the process stepwise.
@@
where $\mathcal{M}_{i}=\bigcup_{j=1}^{i-1}\mathcal{G}_{j}$ represents all the matched subgraphs parsed from the first $i$ subqueries, $\mathcal{G}_{j}$ denotes the maximal matched subgraph in the reference query that aligns with the $i$-th subquery $\hat{q}_{i}$. This formulation prevents repeated rewards for the same reference node and ensures that the overall signal reflects the incremental semantic coverage of the target reference query. Stepwise supervision improves training performance by providing richer intermediate feedback, facilitating the generation of correct SQL queries.
@@ 5 Experimental Setup
Our experiments are primarily conducted on the Spider and BIRD benchmarks. The Spider dataset (Yu et al., 2018) contains 10,181 natural language questions paired with 5,693 complex SQL queries across 138 domains. The BIRD dataset (Li et al., 2024b) consists of 12,751 questions spanning more than 37 professional fields. We used the training split of the Spider dataset for training and the development splits of Spider and BIRD for evaluation. Additionally, we used a subset of the 200k-Text2SQL dataset [^2] https://huggingface.co/datasets/philikai/200k-Text2SQL in a warm-up phase prior to RL training. Further details about the datasets are provided in Appendix A.
@@
We benchmark against representative state-of-the-art (SOTA) Text-to-SQL pipelines spanning in-context learning, self-correction, post-training, and multi-agent paradigms (Pourreza and Rafiei, 2023; Gao et al., 2024a; Pourreza and Rafiei, 2024; Gao et al., 2024b; Pourreza et al., 2025a, b; Zhang et al., 2025). We further compare three widely adopted reward model designs in RL training to validate the effectiveness of the proposed reward models: (i) the execution-based reward EX, widely used in recent studies (Nguyen et al., 2025; Berdnyk and Collery, 2025; Ma et al., 2025; Pourreza et al., 2025b); (ii) an LLM-based Bradley–Terry reward model (BTRM) (Christiano et al., 2017; Zeng et al., 2025b), trained using the DeepSeek-Coder-1.3B-Ins as the backbone, as detailed in Appendix M, to evaluate the efficacy of model-based reward mechanisms; and (iii) AstPM and RelPM (Zhan et al., 2025), which, motivated by recent works (Shojaee et al., 2023; Nguyen et al., 2025; Pourreza et al., 2025b), are incorporated as matching-based reward model baselines.
@@
Consistent with prior work (Pourreza et al., 2025a, b; Zhang et al., 2025), we adopt Execution Accuracy (EX) as the primary metric for comparisons with state-of-the-art systems. In addition, we report Test-Suite Accuracy (TS), which evaluates predicted queries against suites of distilled or fuzzed database variants under varied conditions (Zhong et al., 2020). Since TS mitigates the false positives that may arise from insufficiently discriminative databases, it provides a more robust estimate of semantic correctness (Gao et al., 2024a; Li et al., 2024a; Yang et al., 2024). We therefore use TS to more accurately quantify the gains achieved by our reward model.
@@
All experiments are conducted on several edge servers. We extend the verl [^3] https://github.com/volcengine/verl and DeepSpeed-Chat [^4] https://github.com/deepspeedai/DeepSpeed (Yao et al., 2023) framework to support a comparison of multiple reward models, including execution-based, LLM-based, and matching-based rewards.
Prior to RL training, we performed SFT using two cold-start datasets. First, we sampled a subset from the 200k-Text2SQL dataset, matching the size of the Spider training set, and trained the base model for two epochs. To promote the generation of CTE SQL queries in the stepwise reward PPO experiments, we converted BIRD data into CTE format to prepare a warm-up dataset referred to as CTE-SFT. During supervised fine-tuning (SFT), we employ the AdamW optimizer (initial learning rate 5e-6, batch size 4) and use a polynomial scheduler to decay the learning rate throughout training. Additional details about hyperparameter of SFT and RL are provided in Appendix C.
@@ 6 Results
Method Method Detail Base Model Spider EX BIRD EX Avg.
DIN-SQL (Pourreza and Rafiei, 2023) In-context learning GPT-4 82.88 50.72 66.80
DAIL-SQL (Gao et al., 2024a) In-context learning GPT-4 74.47 54.76 64.62
DTS-SQL (Pourreza and Rafiei, 2024) Schema-linking, SFT DeepSeek-7B - 55.80 -
XiYan-SQL (Gao et al., 2024b) Multi-Agent QwenCoder-32B - 67.01 -
CHASE-SQL (Pourreza et al., 2025a) Multi-Agent Gemini-1.5-pro - 74.46 -
Reasoning-SQL (Pourreza et al., 2025b) Multi-Agent, GRPO Qwen2.5-Coder-7B 78.72 64.01 71.37
Reward-SQL (Zhang et al., 2025) GRPO Qwen2.5-Coder-7B 77.08 59.70 68.39
Graph-Reward-SQL (Ours) GRPO Qwen2.5-Coder-7B 81.62 63.04 72.33
Table 2: Performance comparison of Graph-Reward-SQL and baseline models on the Spider and BIRD dataset. Baseline results are reported from their original publications.
We benchmark our method against recent SOTA Text-to-SQL systems, including in-context learning and multi-agent approaches. As shown in Table 2, Graph-Reward-SQL reaches 81.62% EX on Spider and 63.04% on BIRD trained with GRPO on the Qwen2.5-Coder-7B-Ins model. Our reward model framework delivers performance competitive with the Reasoning-SQL, even as the latter employs a multi-agent system and five reward models that incorporate both EX and LLM-based judgment.
@@ 6.1 Reward Performance Comparison
GMNScore can replace the EX, thereby eliminating dependence on SQL execution and database environments. As demonstrated in Table 3, GMNScore achieves the highest average TS for the 1.3B and 6.7B models, highlighting the importance of well-designed reward signals in RL. Another notable observation is that RelPM outperforms AstPM, with improvements of 2.53% and 1.71% for the two model sizes, respectively. The better performance of the former can be attributed to the use of normalized logical plans for SQL parsing in ROT, which are less susceptible to surface-level syntactic differences. ROT also provides an effective representation way for our proposed GMNScore and StepRTM reward models. GMNScore learns deep semantic information via graph-level embeddings, bypassing the need for execution-result comparisons and thus mitigating false-positive noise. Additionally, GMNScore eliminates the necessity of constructing and maintaining databases, offering a lightweight solution for large-scale RL-based Text-to-SQL. Case studies are provided in Appendix  S.
@@
DeepSeek-Coder-1.3B-Ins + PPO 39.56 11.34 25.45
RelPM w/ SFT 62.86 14.67 38.77
RelPM w/ CTE-SFT 62.57 19.36 40.97
RelPM & StepRTM w/ CTE-SFT 64.31 19.43 41.87
EX w/ SFT 65.28 17.21 41.25
EX w/ CTE-SFT 65.09 18.97 42.03
EX & StepRTM w/ CTE-SFT 67.89 19.75 43.82
GMNScore w/ SFT 67.70 16.10 41.90
GMNScore w/ CTE-SFT 68.57 20.40 44.49
GMNScore & StepRTM w/ CTE-SFT 68.67 21.97 45.32
Table 4: TS performance of the DeepSeek-Coder-1.3B-Ins model trained with StepRTM integration. Both SFT and CTE-SFT refer to different warm-up datasets.
The integration of StepRTM as a stepwise reward further enhances performance. As shown in Table 4, combining CTE-SFT with StepRTM consistently results in metric improvements across various outcome reward models. Notably, our framework, which integrates GMNScore alongside StepRTM, achieves the highest result. Specifically, we observe a 5.87% improvement on the BIRD dataset and a 0.97% increase on the Spider dataset. These observations suggest that the BIRD dataset, which is inherently more challenging due to its diverse database and query complexity, benefits more significantly from our proposed stepwise reward.
@@ 6.2 Effectiveness of GMNScore with GRPO
GMNScore performs robustly across both PPO and GRPO protocols. As shown in Figure 3, the results consistently demonstrate that GMNScore outperforms EX in these two RL protocols, underscoring its robustness and effectiveness. We report the average performance across two benchmarks, with detailed results provided in Appendix K.
[…]
EX 1.088s
BTRM 0.095s 1.35B 9304MB
GMNScore 0.023s 3.99M 83MB
Table 5: Average per-sample cost of reward scoring.
6.3 Cost of GMNScore
We compare the cost of different reward models by running one PPO epoch and measuring the average reward-scoring latency per sample. As shown in Table 5, GMNScore is the most efficient. In contrast, EX is the slowest due to repeated database calls. From a model-size perspective, GMNScore is highly compact, with 3.99M parameters, whereas the widely used BTRM is substantially larger at 1.35B parameters. GMNScore is trained once offline and can be reused across multiple RL experiments, which significantly reduces overall computational overhead. Additional details on the measurement setup are provided in Appendix G.
@@ 6.3 Cost of GMNScore
6.4 Intrinsic Evaluation of Reward Models
Reward Model Spider-pair dev BIRD-pair dev
AstPM 82.81% 80.38%
RelPM 84.42% 83.57%
BTRM 89.24% 84.77%
EX 96.37% 92.67%
GMNScore 97.62% 94.14%
Table 7: Area Under the Curve (AUC) of reward models on Spider-dev-pair and BIRD-dev-pair.
Area Under the Curve (AUC) is used as a statistically principled measure of similarity-evaluation accuracy (Zhan et al., 2025); its alignment with RL reward-model objectives makes it a rigorous criterion for validating reward-model performance. As shown in Table 7, GMNScore achieves 97.62% AUC on Spider-dev-pair and 94.14% on BIRD-dev-pair. Equivalence labels in the these two dev-pair sets were manually verified to minimize false positives, improving the reliability of these AUC estimates. Collectively, the results indicate that GMNScore provides accurate and consistent preference judgments, laying a solid foundation for subsequent RL training. Details on training and evaluation datasets are provided in Appendix E.
6.5 Case Study: CTE SQL with StepRTM
Table 6 presents two cases that demonstrate how the stepwise reward model enhances both correctness and readability. Each case compares the reference SQL, a failed SQL query generated by a model trained solely with an outcome reward model, and a CTE SQL query generated by a model trained with StepRTM. In the first case, the failed SQL incorrectly retrieves data from the ‘comments’ table instead of the intended ‘posts’ table. The CTE SQL resolves this by decomposing the task into clear subqueries: first locating the target user, then aggregating the scores of that user’s posts. In the second case, the failed SQL hard-codes the gender identifier as ’1’, leading to errors in filtering. In contrast, the CTE SQL uses two dedicated subqueries to correctly filter male superheroes, extract their superpowers, finally combine the results.
@@ 7 Discussion
The GMNScore introduced in this paper offers an alternative to EX while remaining fully compatible with other reward models. As detailed in Appendix I, we extend our investigation beyond the StepRTM integration (Section 6.1) by applying hybrid outcome reward models, which further improve performance. This finding is consistent with previous work using multiple outcome rewards (Pourreza et al., 2025b; Ma et al., 2025).
@@ 8 Conclusion
We propose Graph-Reward-SQL, a reward model framework for RL-based Text-to-SQL that replaces execution-based rewards with the GMNScore outcome reward model and the StepRTM stepwise reward model. By eliminating the dependency on execution during training, it significantly improves training efficiency. Extensive experiments demonstrate that our proposed reward models achieve superior performance. The Graph-Reward-SQL framework establishes a new direction toward scalable and efficient RL-based Text-to-SQL tasks.